OpenStudy (anonymous):

Find the general solution to the first-order differential equation 2x(y+1) - yy' = 0 will post choices

OpenStudy (anonymous):

\[A) 2x ^{2}-y ^{2}\ln \left| y+1 \right|=C\] \[B) x ^{2}(y+1)^{2}-y ^{2}=C\] \[C)\ln \left| y+1 \right|+x ^{2}-y= C\] \[D) x ^{2}=y+(1/(y+1)^{2})+C\]

OpenStudy (anonymous):

@crzymom2u How?

OpenStudy (anonymous):

@crzymom2u how?

OpenStudy (nikvist):